If the question is x+23=8x-5 what is the answer?
Answer by ikleyn(52786)

.

Hello,


    x + 23 = 8x - 5


is not a question. It is an equation.


A question is how to solve it . . . 



                SOLUTION


    x + 23 = 8x - 5


Move x from left side to right side, changing its sign.

    23     = 8x - 5 - x


Combine like terms on the left side

    23     = 7x - 5


Move (-5) from right side to the left side, changing the sign

    23 + 5 = 7x


Combine like terms in the left side

   28      = 7x


Divide both sides by 7

    x      = 28/7 = 4.


ANSWER.  The solution is  x= 4.

Solved.

----------------

If you want to learn on how to solve simple and simplest linear equations,  read the lesson
    - HOW TO solve a linear equation
in this site.

Consider this lesson as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.

Happy learning  (!)